OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (–1, 5) and has a slope of –4? y = _____ x + _____

OpenStudy (tanner23456):

Use this formula y-y1 = m(x-x1) x1 = first coordinate y1= second coordinate m = slope